OpenStudy (aaronandyson):

A body is thrown up with a velocity of 40ms^-1 .At the same time another body is dropped from a height of 40m.Their relative acceleration after 1.3s is?

OpenStudy (faiqraees):

Can you confirm the choices here

OpenStudy (aaronandyson):

Options: 4g g/2 2g 0

OpenStudy (faiqraees):

Do you want to venture a guess?

OpenStudy (aaronandyson):

Like what?

OpenStudy (faiqraees):

Do you have an opinion about what the answer is going to be?

OpenStudy (aaronandyson):

I have no idea...

OpenStudy (faiqraees):

Okay so lets call the up going body A and falling body B What are the acceleration (relative to a stationary observer) of both the bodies?

OpenStudy (aaronandyson):

a of A to the a of B?

OpenStudy (faiqraees):

no just the acceleration of A and the acceleration of B? (Not relative to each other)

OpenStudy (aaronandyson):

v-u/t = a

OpenStudy (faiqraees):

What is the acceleration of a free-falling body on Earth (relative to ground)? Note:- Dont read the brackets, if you cant understand what it means. Just answer the non-bracket part.

OpenStudy (aaronandyson):

9.81

OpenStudy (faiqraees):

Yes, or in other words g. So what is the acceleration of both the bodies?(Relative to the ground)

OpenStudy (aaronandyson):

both should be g ..?

OpenStudy (faiqraees):

Perfect. What is the direction of their acceleration? Away from ground or towards the ground?

OpenStudy (aaronandyson):

A is upwards and B is downwards.

OpenStudy (faiqraees):

A bit incorrect. Remember, velocity and acceleration are different. You can go up and have an acceleration of downwards, in other words decreasing velocity.

OpenStudy (aaronandyson):

vel of a increasing and that of b is decreasing?

OpenStudy (faiqraees):

Okay a quick fact. Acceleration of a body in Earth is always directed towards the ground. No matter which direction you're going. You will always have a component of acceleration downwards (or in complex way, towards the centre of Earth) So can you again make a informed decision about the directions of their acceleration

OpenStudy (aaronandyson):

Both are directed towards centre of the earth.

OpenStudy (faiqraees):

Okay so both are accelerating in the same direction at the same rate. Thus their relative acceleration will be?

OpenStudy (faiqraees):

Okay consider a case. You are sitting in a train and watching another person on an another train which is going at the same speed in the same direction as your train. You will consider person (his relative velocity to you) A. Stationary B. Same as train's speed C. Double the train's speed?

OpenStudy (aaronandyson):

A?

OpenStudy (faiqraees):

Yes. Same is the case with acceleration. Same direction and same magnitude of acceleration will result in relative acceleration of a magnitude?

OpenStudy (aaronandyson):

wait.. a is -g right?

OpenStudy (aaronandyson):

-g+g = 0

OpenStudy (faiqraees):

Yes Correct

OpenStudy (faiqraees):

The correct solution is Accel of A - Accel of B = Accel of A relative to B (Since they both are in same direction) Accel of A relative to B = -g - (-g)=-g+g=0